How do I put the following equations in point slope form, and how do I have to substitute to find the answer of 5x+6y=-11 and 3x+y=-4 ?
Answer by rfer(16322)

3x+y=-4
y=-3x-4
------------
5x+6y=-11
5x+6(-3x-4)=-11
5x-18x-24=-11
-13x=-11+24
-13x=13
x=13/-13
x=-1
--------------
5*-1+6y=-11
-5+6y=-11
6y=-11+5
6y=-6
y=-6/6
y=-1
------------
3*-1-1=-4
-3-1=-4
You do the rest